# Temporal Valued Constraint Satisfaction Problems

### v1

@@ 2.2. Valued constraint satisfaction problems
Let $\tau$ be a relational signature. An atomic $\tau$-expression is an expression of the form $R(x_{1},\dots,x_{k})$ for $R\in\tau$, $(=)_{0}^{\infty}(x_{1},x_{2})$, or $(\emptyset)_{0}^{\infty}(x_{1})$ where $x_{1},\dots,x_{k}$ are (not necessarily distinct) variable symbols. A $\tau$-expression is an expression $\phi$ of the form $\sum_{i\leq m}\phi_{i}$ where $m\in{\mathbb{N}}$ and $\phi_{i}$ for $i\in\{1,\dots,m\}$ is an atomic $\tau$-expression. Note that the same atomic $\tau$-expression might appear several times in the sum. We write $\phi(x_{1},\dots,x_{n})$ for a $\tau$-expression where all the variables are from the set $\{x_{1},\dots,x_{n}\}$. If ${\mathfrak{A}}$ is a valued $\tau$-structure, then a $\tau$-expression $\phi(x_{1},\dots,x_{n})$ defines over ${\mathfrak{A}}$ a member of ${\mathscr{R}}_{A}^{(n)}$, which we denote by $\phi^{{\mathfrak{A}}}$. If $\phi$ is the empty sum then $\phi^{{\mathfrak{A}}}$ is constant 00.
@@
Let ${\mathfrak{A}}$ be a valued structure over a finite signature $\tau$. The valued constraint satisfaction problem for ${\mathfrak{A}}$, denoted by $\operatorname{VCSP}({\mathfrak{A}})$, is the computational problem to decide for a given $\tau$-expression $\phi(x_{1},\dots,x_{n})$ and a given $u\in{\mathbb{Q}}$ whether there exists $t\in A^{n}$ such that $\phi^{{\mathfrak{A}}}(t)\leq u$. We refer to $\phi$ as an instance of $\operatorname{VCSP}({\mathfrak{A}})$, and to $u$ as the threshold. We also refer to the pair $(\phi,u)$ as a (positive or negative) instance of $\operatorname{VCSP}({\mathfrak{A}})$. Tuples $t\in A^{n}$ such that $\phi^{{\mathfrak{A}}}(t)\leq u$ are called a solution for $(\phi,u)$. The cost of $\phi$ (with respect to ${\mathfrak{A}}$) is defined to be
$\inf_{t\in A^{n}}\phi^{{\mathfrak{A}}}(t).$
In some contexts, it will be beneficial to consider only a given $\tau$-expression $\phi$ to be the input of $\operatorname{VCSP}({\mathfrak{A}})$ (rather than $\phi$ and the threshold $u$) and a tuple $t\in A^{n}$ will then be called a solution for $\phi$ if the cost of $\phi$ equals $\phi^{{\mathfrak{A}}}(t)$. Note that in general there might not be any solution. If there exists a tuple $t\in A^{n}$ such that $\phi^{{\mathfrak{A}}}(t)<\infty$ then $\phi$ is called satisfiable.
Example 2.2.
Let $\tau=\{E\}$ and let ${\mathfrak{A}}$ be a valued $\tau$-structure on the domain $\mathbb{Q}$ where $E^{{\mathfrak{A}}}=(<)_{0}^{1}$. Then every $\tau$-expression can be interpreted as a (not necessarily simple) digraph with the edge relation $E$ and every digraph corresponds to a $\tau$-expression. The cost of every instance of $\operatorname{VCSP}({\mathfrak{A}})$ is equal to the number of edges that have to be removed to make the digraph acyclic. Therefore, $\operatorname{VCSP}({\mathfrak{A}})$ is the minimum feedback arc set problem.
Remark 2.3.
@@
If ${\mathfrak{A}}$ be a relational $\tau$-structure, then $\operatorname{CSP}({\mathfrak{A}})$ is the problem of deciding satisfiability of conjunctions of atomic $\tau$-formulas in ${\mathfrak{A}}$. Note that every $\tau$-expression $\phi(x_{1},\dots,x_{k})$ defines a crisp relation and can be viewed as an existentially quantified conjunction of atomic formulas, i.e., primitive positive formula, which defines the same relation. Therefore, $\operatorname{VCSP}({\mathfrak{A}})$ and $\operatorname{CSP}({\mathfrak{A}})$ are essentially the same problem.
2.3. Automorphisms
@@
Let $k\in{\mathbb{N}}$, let $R\in{\mathscr{R}}^{(k)}_{A}$, and let $\alpha$ be a permutation of $A$. Then $\alpha$ preserves $R$ if for all $t\in A^{k}$ we have $R(\alpha(t))=R(t)$. If ${\mathfrak{A}}$ is a valued structure with domain $A$, then an automorphism of ${\mathfrak{A}}$ is a permutation of $A$ that preserves all valued relations of $R$. The set of all automorphisms of ${\mathfrak{A}}$ is denoted by $\operatorname{Aut}({\mathfrak{A}})$, and forms a group with respect to composition. If ${\mathfrak{B}}$ is a valued structure and we write $\operatorname{Aut}({\mathfrak{B}})\subseteq\operatorname{Aut}({\mathfrak{A}})$ or $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}({\mathfrak{A}})$, we implictly assume that ${\mathfrak{A}}$ and ${\mathfrak{B}}$ have the same domain.
@@
Let $k\in{\mathbb{N}}$. An orbit of $k$-tuples of a permutation group $G$ on a set $A$ is a set of the form $\{\alpha(t)\mid\alpha\in G\}$ for some $t\in A^{k}$. A permutation group $G$ on a countable set is called oligomorphic if for every $k\in{\mathbb{N}}$ there are finitely many orbits of $k$-tuples in $G$ [18]. For example, $\operatorname{Aut}(\mathbb{Q};<)$ and therefore every permutation group on $\mathbb{Q}$ that contains $\operatorname{Aut}(\mathbb{Q};<)$ is oligomorphic. If ${\mathfrak{A}}$ is a relational structure with an oligomorphic automorphism group and $R\subseteq A^{k}$, then $R$ is first-order definable over ${\mathfrak{A}}$ if and only if $R$ is preserved by $\operatorname{Aut}({\mathfrak{A}})$, see, e.g., [3, Theorem 4.2.9].
@@
Let ${\mathfrak{A}}$ be a valued $\tau$-structure and ${\mathfrak{B}}$ a relational structure. Suppose that $\operatorname{Aut}({\mathfrak{B}})$ is oligomorphic and $\operatorname{Aut}({\mathfrak{B}})\subseteq\operatorname{Aut}({\mathfrak{A}})$ (and hence $\operatorname{Aut}({\mathfrak{A}})$ is oligomorphic). Let $R\in\tau$ be of arity $k$. Then $R^{{\mathfrak{A}}}$ attains only finitely many values by the oligomorphicity of $\operatorname{Aut}({\mathfrak{A}})$. Moreover, if for some $s,t\in A^{k}$ we have $R^{{\mathfrak{A}}}(s)\neq R^{{\mathfrak{A}}}(t)$, then $s$ and $t$ lie in a different orbit of $\operatorname{Aut}({\mathfrak{B}})$. Therefore, for every value $a\in\mathbb{Q}\cup\{\infty\}$, there is a union $U_{a}$ of orbits of $k$-tuples under the action of $\operatorname{Aut}({\mathfrak{B}})$ such that $R^{{\mathfrak{A}}}(t)=a$ if and only if $t\in U_{a}$. Since $U_{a}$ is preserved by $\operatorname{Aut}({\mathfrak{B}})$, it is first-order definable over ${\mathfrak{B}}$ by a formula $\phi_{a}$. Hence, $R$ can be given by a list of values $a$ in the range of $R$ and first-order formulas $\phi_{a}$ over ${\mathfrak{B}}$. Such a collection
$((R,a,\phi_{a})\mid R\in\tau,\exists t\in A^{k}(R(t)=a))$
will be called a first-order definition of ${\mathfrak{A}}$ in ${\mathfrak{B}}$. Clearly, if a valued structure ${\mathfrak{A}}$ has a first-order definition in a relational structure ${\mathfrak{B}}$, then $\operatorname{Aut}({\mathfrak{B}})\subseteq\operatorname{Aut}({\mathfrak{A}})$. Note that for some structures ${\mathfrak{B}}$ such as $(\mathbb{Q};=)$ and $(\mathbb{Q};<)$, the formulas $\phi_{a}$ can be chosen to be quantifier-free, and hence as disjunctions of conjunctions of atomic formulas over ${\mathfrak{B}}$ (in fact, this is the case for every homogeneous structure with a finite relational signature). We will use first-order definitions of valued structures to be able to give valued structures as an input to decision problems (see Remark 4.7 and Proposition 5.25).
2.4. Expressive power
@@
  • projecting if $R^{\prime}$ is of arity $k$, $R$ is of arity $k+n$ and for all $s\in A^{k}$
    $R^{\prime}(s)=\inf_{t\in A^{n}}R(s,t).$
@@
Note that $\inf_{t\in A^{n}}R(s,t)$ in item (1) might be irrational or $\infty$- ∞. If this is the case, then $\inf_{t\in A^{n}}R(s,t)$ does not express a valued relation because valued relations must have weights from ${}{\mathbb{Q}}\cup\{\infty\}$. However, if $R$ is preserved by all permutations of an oligomorphic automorphism group, then $R$ attains only finitely many values and therefore this is never the case.
If $S\subseteq\mathscr{R}_{A}$, then an atomic expression over $S$ is an atomic $\tau$-expression where $\tau=S$. We say that $S$ is closed under forming sums of atomic expressions if it contains all valued relations defined by sums of atomic expressions over $S$.
Definition 2.5 (valued relational clone).
A valued relational clone (over $A$) is a subset of ${\mathscr{R}}_{A}$ that is closed under forming sums of atomic expressions, projecting, shifting, non-negative scaling, $\operatorname{Feas}$, and $\operatorname{Opt}$. For a valued structure ${\mathfrak{A}}$ with the domain $A$, we write $\langle{\mathfrak{A}}\rangle$ for the smallest valued relational clone that contains the valued relations of ${\mathfrak{A}}$. If $R\in\langle{\mathfrak{A}}\rangle$, we say that ${\mathfrak{A}}$ expresses $R$.
Remark 2.6.
Note that if a valued relational clone $\mathscr{C}$ contains a set $S\subseteq{\mathscr{R}}_{A}$ of crisp relations, then every relation which is primitively positively definable from $S$ is in $\mathscr{C}$ by forming a sum of the corresponding atomic expressions and projecting on the variables that are not existentially quantified. Therefore, valued relational clones are a generalization of relational clones. Moreover, if ${\mathfrak{A}}$ is a relational structure and $R\in\langle{\mathfrak{A}}\rangle$, then $R$ is essentially crisp and $\operatorname{Feas}(R)$ is primitively positively definable from ${\mathfrak{A}}$; this is easily verified by induction.
The following lemma is the main motivation for the concept of expressibility.
Lemma 2.7 (Lemma 4.6 in [15]).
@@
We now introduce notation that enables us to talk about the crisp relations expressible in a valued structure, which turn out to be essential to understanding temporal VCSPs.
Definition 2.8.
Let ${\mathfrak{A}}$ be a valued structure. Then $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ denotes the set of valued relations
$\{R\in\langle{\mathfrak{A}}\rangle\mid R\text{ of arity }k,\forall a~{}\in A^{k}\colon R(a)\in\{0,\infty\}\}.$
In words, $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ contains all crisp relations that can be expressed in ${\mathfrak{A}}$. We finish this section with a remark on expressibility in essentially crisp valued structures.
Remark 2.9.
Let ${\mathfrak{A}}$ be an essentially crisp valued $\tau$-structure. For every $R\in\tau$, let $a_{R}\in\mathbb{Q}$ be such that $R^{{\mathfrak{A}}}$ only attains values in $\{a_{R},\infty\}$; such an $a_{R}$ exists because ${\mathfrak{A}}$ is essentially crisp. Then $R^{{\mathfrak{A}}}=\operatorname{Feas}(R^{{\mathfrak{A}}})+a_{R}$. Therefore, $\langle{\mathfrak{A}}\rangle=\langle\operatorname{Feas}({\mathfrak{A}})\rangle$ and, by Remark 2.6, $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ consists of precisely those relations that are primitively positively definable in $\operatorname{Feas}({\mathfrak{A}})$. By Lemma 2.7 and Remark 2.3, there is a polynomial-time reduction from $\operatorname{VCSP}({\mathfrak{A}})$ to $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ and vice versa.
2.5. Pp-constructions
Next, we introduce a concept of pp-constructions which give rise to polynomial-time reductions between VCSPs. The acronym ‘pp’ stands for primitive positive, since the concept of pp-constructions for relational structures is a generalization of primitive positive definitions used for reductions between CSPs.
Definition 2.10 (pp-power).

### v2

@@ 2.2 Valued constraint satisfaction problems
Let $\tau$ be a relational signature. An atomic $\tau$-expression is an expression of the form $R(x_{1},\dots,x_{k})$ for $R\in\tau$, $(=)_{0}^{\infty}(x_{1},x_{2})$, or $\bot(x_{1})$ where $x_{1},\dots,x_{k}$ are (not necessarily distinct) variable symbols. A $\tau$-expression is an expression $\phi$ of the form $\sum_{i\leq m}\phi_{i}$ where $m\in{\mathbb{N}}$ and $\phi_{i}$ for $i\in\{1,\dots,m\}$ is an atomic $\tau$-expression. Note that the same atomic $\tau$-expression might appear several times in the sum. We write $\phi(x_{1},\dots,x_{n})$ for a $\tau$-expression where all the variables are from the set $\{x_{1},\dots,x_{n}\}$. If $\mathfrak{A}$} is a valued $\tau$-structure, then a $\tau$-expression $\phi(x_{1},\dots,x_{n})$ defines over $\mathfrak{A}$} a member of ${\mathscr{R}}_{A}^{(n)}$ in the usual way, which we denote by $\phi^{{\mathfrak{A}}}$ (for example, if $\phi(x,y)=R(x)+S(y)$ and $\mathfrak{A}$} is an $\{R,S\}$-structure, then $\phi^{\mathfrak{A}}(x,y)=R^{\mathfrak{A}}(x)+S^{\mathfrak{A}}(y)$ for all $x,y\in A$). If $\phi$ is the empty sum, then $\phi^{{\mathfrak{A}}}$ is constant 0.
@@
Let $\mathfrak{A}$} be a valued structure over a finite signature $\tau$. The valued constraint satisfaction problem for $\mathfrak{A}$}, denoted by $\operatorname{VCSP}({\mathfrak{A}})$, is the computational problem to decide for a given $\tau$-expression $\phi(x_{1},\dots,x_{n})$ and a given $u\in{\mathbb{Q}}$ whether there exists $t\in A^{n}$ such that $\phi^{{\mathfrak{A}}}(t)\leq u$. We refer to $\phi$ as an instance of $\operatorname{VCSP}({\mathfrak{A}})$, and to $u$ as the threshold. We also refer to the pair $(\phi,u)$ as a (positive or negative) instance of $\operatorname{VCSP}({\mathfrak{A}})$. A tuple $t\in A^{n}$ such that $\phi^{{\mathfrak{A}}}(t)\leq u$ is called a solution for $(\phi,u)$. The cost of $\phi$ (with respect to $\mathfrak{A}$}) is defined to be $\inf_{t\in A^{n}}\phi^{{\mathfrak{A}}}(t).$ In some contexts, it will be beneficial to consider only a given $\tau$-expression $\phi$ to be the input of $\operatorname{VCSP}({\mathfrak{A}})$ (rather than $\phi$ and the threshold $u$) and a tuple $t\in A^{n}$ will then be called a solution for $\phi$ if the cost of $\phi$ equals $\phi^{{\mathfrak{A}}}(t)$. In general, there might not be any solution. If there exists $t\in A^{n}$ such that $\phi^{{\mathfrak{A}}}(t)<\infty$ then $\phi$ is called satisfiable. To give an example of a VCSP, note that $\operatorname{VCSP}(\mathbb{Q};(<)_{0}^{1})$ is the minimum feedback arc set problem for directed multigraphs.
@@
If $\mathfrak{A}$} is a relational $\tau$-structure, then $\operatorname{CSP}({\mathfrak{A}})$ is the problem of deciding satisfiability of conjunctions of atomic $\tau$-formulas in $\mathfrak{A}$}. Therefore, if $\mathfrak{A}$} is a relational structure then $\operatorname{VCSP}({\mathfrak{A}})$ and $\operatorname{CSP}({\mathfrak{A}})$ are essentially the same problem.
2.3 Automorphisms
@@
Let $k\in{\mathbb{N}}$, let $R\in{\mathscr{R}}^{(k)}_{A}$, and let $\alpha$ be a permutation of $A$. Then $\alpha$ preserves $R$ if for all $t\in A^{k}$ we have $R(\alpha(t))=R(t)$. If $\mathfrak{A}$} is a valued structure with domain $A$, then an automorphism of $\mathfrak{A}$} is a permutation of $A$ that preserves all valued relations of $R$. The set of all automorphisms of $\mathfrak{A}$} is denoted by $\operatorname{Aut}({\mathfrak{A}})$, and forms a group with respect to composition. If $\mathfrak{B}$} is a valued structure and we write $\operatorname{Aut}({\mathfrak{B}})\subseteq\operatorname{Aut}({\mathfrak{A}})$ or $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}({\mathfrak{A}})$, we implicitly assume that $\mathfrak{A}$} and $\mathfrak{B}$} have the same domain.
@@
Let $k\in{\mathbb{N}}$. An orbit of $k$-tuples of a permutation group $G$ on a set $A$ is a set of the form $\{\alpha(t)\mid\alpha\in G\}$ for some $t\in A^{k}$. A permutation group $G$ on a countable set is called oligomorphic if for every $k\in{\mathbb{N}}$ there are finitely many orbits of $k$-tuples in $G$ [20]. For example, $\operatorname{Aut}(\mathbb{Q};<)$ and therefore every permutation group on $\mathbb{Q}$ that contains $\operatorname{Aut}(\mathbb{Q};<)$ is oligomorphic. If $\mathfrak{A}$} is a relational structure with an oligomorphic automorphism group and $R\subseteq A^{k}$, then $R$ is first-order definable over $\mathfrak{A}$} if and only if $R$ is preserved by $\operatorname{Aut}({\mathfrak{A}})$ (this theorem, including the definition of first-order logic, is treated, e.g., in [17] (Theorem 4.2.9)).
@@
Let $\mathfrak{A}$} be a valued $\tau$-structure and $\mathfrak{B}$} a relational structure. Suppose that $\operatorname{Aut}({\mathfrak{B}})$ is oligomorphic and $\operatorname{Aut}({\mathfrak{B}})\subseteq\operatorname{Aut}({\mathfrak{A}})$ (and hence $\operatorname{Aut}({\mathfrak{A}})$ is oligomorphic). Let $R\in\tau$ be of arity $k$. Then $R^{{\mathfrak{A}}}$ attains only finitely many values by the oligomorphicity of $\operatorname{Aut}({\mathfrak{A}})$. Moreover, if for some $s,t\in A^{k}$ we have $R^{{\mathfrak{A}}}(s)\neq R^{{\mathfrak{A}}}(t)$, then $s$ and $t$ lie in a different orbit of $\operatorname{Aut}({\mathfrak{B}})$. Therefore, for every value $a\in\mathbb{Q}\cup\{\infty\}$, there is a union $U_{a}$ of orbits of $k$-tuples under the action of $\operatorname{Aut}({\mathfrak{B}})$ such that $R^{{\mathfrak{A}}}(t)=a$ if and only if $t\in U_{a}$. Since $U_{a}$ is preserved by $\operatorname{Aut}({\mathfrak{B}})$, it is first-order definable over $\mathfrak{B}$} by a formula $\phi_{a}$. Hence, $R$ can be given by a list of values $a$ in the range of $R$ and first-order formulas $\phi_{a}$ over $\mathfrak{B}$}. Such a collection $((R,a,\phi_{a})\mid R\in\tau,\exists t\in A^{k}(R(t)=a))$ will be called a first-order definition of $\mathfrak{A}$} in $\mathfrak{B}$}. Clearly, if a valued structure $\mathfrak{A}$} has a first-order definition in a relational structure $\mathfrak{B}$}, then $\operatorname{Aut}({\mathfrak{B}})\subseteq\operatorname{Aut}({\mathfrak{A}})$. Note that for some structures $\mathfrak{B}$} such as $(\mathbb{Q};=)$ and $(\mathbb{Q};<)$, the formulas $\phi_{a}$ can be chosen to be quantifier-free, and hence as disjunctions of conjunctions of atomic formulas over $\mathfrak{B}$} (in fact, this is the case for every homogeneous structure with a finite relational signature). We will use first-order definitions of valued structures to be able to give valued structures as an input to decision problems (see Proposition 56).
2.4 Expressive power
@@
  1. 1.
    projecting if $R^{\prime}$ is of arity $k$, $R$ is of arity $k+n$ and for all $s\in A^{k}$, $R^{\prime}(s)=\inf_{t\in A^{n}}R(s,t)$.
  2. 2.
@@
Note that $\inf_{t\in A^{n}}R(s,t)$ in item 1 might be irrational or -$\infty$. If this is the case, then $\inf_{t\in A^{n}}R(s,t)$ does not express a valued relation because valued relations must have weights from $\mathbb{Q}$$\cup\{\infty\}$. However, if $R$ is preserved by all permutations of an oligomorphic automorphism group, then $R$ attains only finitely many values and therefore this is never the case.
If $S\subseteq\mathscr{R}_{A}$, then an atomic expression over $S$ is an atomic $\tau$-expression where $\tau=S$. We say that $S$ is closed under forming sums of atomic expressions if for every $n\in\mathbb{N}$, $R_{1},\dots,R_{n}\in S$ of arity $k_{1},\dots,k_{n}$, respectively, and $i_{1}^{j},\dots,i_{k_{j}}^{j}\in[k_{j}]$ for $j\in\{1,\dots,n\}$, the set $S$ also contains the valued relation $R$ of arity $k$ defined by
$R(x_{1},\dots,x_{n}):=\sum_{j=1}^{n}R_{j}(x_{i_{1}^{j}},\dots,x_{i_{k_{j}}^{j}}).$
Definition 3 (valued relational clone)
A valued relational clone (over $A$) is a subset of ${\mathscr{R}}_{A}$ that is closed under forming sums of atomic expressions, projecting, shifting, non-negative scaling, $\operatorname{Feas}$, and $\operatorname{Opt}$; we will refer to expressions formed this way as pp-expressions. For a valued structure $\mathfrak{A}$} with the domain $A$, we write $\langle{\mathfrak{A}}\rangle$ for the smallest valued relational clone that contains the valued relations of $\mathfrak{A}$}. If $R\in\langle{\mathfrak{A}}\rangle$, we say that $\mathfrak{A}$} $p$-expresses $R$.
Note that every relation which is primitively positively definable from a set $S\subseteq{\mathscr{R}}_{A}$ of crisp relations lies in $\langle S\rangle$. Moreover, if $\mathfrak{A}$} is a relational structure and $R\in\langle{\mathfrak{A}}\rangle$, then $R$ is essentially crisp and $\operatorname{Feas}(R)$ is primitively positively definable from $\mathfrak{A}$}; this is easily verified by induction. This justifies the use of the acronym ‘pp’ that stands for primitive positive, because pp-expressions generalize primitive positive definitions.
The following lemma is the main motivation for the concept of pp-expressibility.
Lemma 4 (Lemma 4.6 in [16])
@@
We now introduce notation that enables us to talk about the crisp relations pp-expressible in a valued structure, which turn out to be essential to understanding temporal VCSPs.
Definition 5
Let $\mathfrak{A}$} be a valued structure. Then $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ denotes the set of valued relations $\{R\in\langle{\mathfrak{A}}\rangle\mid R\text{ of arity }k,\forall a~\in A^{k}\colon R(a)\in\{0,\infty\}\}.$
In words, $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ contains all crisp relations that can be pp-expressed in $\mathfrak{A}$}. If $\mathfrak{A}$} is essentially crisp, then $\langle{\mathfrak{A}}\rangle=\langle\operatorname{Feas}({\mathfrak{A}})\rangle$ and $\langle{\mathfrak{A}}\rangle_{0}^{\infty}$ consists of precisely those relations that are primitively positively definable in $\operatorname{Feas}({\mathfrak{A}})$. In this case we obtain a polynomial-time reduction from $\operatorname{VCSP}({\mathfrak{A}})$ to $\operatorname{CSP}(\operatorname{Feas}({\mathfrak{A}}))$ and vice versa by Lemma 4.
2.5 Pp-constructions
Next, we introduce a concept of pp-constructions which give rise to polynomial-time reductions between VCSPs.
Definition 6 (pp-power)